OpenStudy (anonymous):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. The value of g(-2) is smaller than the value of g(4). The value of g(-2) is the same as the value of g(4). The values of g(-2) and g(4) cannot be compared. The value of g(-2) is larger than the value of g(4).

OpenStudy (kewlgeek555):

Okay, so what you have to do it plug in the g(-2) and solve, and then plug in the g(4) and solve. Let me do the g(-2) and you can do the g(4). \[g(-2)=-3(-2)+4\]\[g(-2)=-6+4\]\[g(-2)=-2\] Are you comfortable with doing g(4) by yourself?

OpenStudy (anonymous):

g(4)=-8 is what I got?

OpenStudy (anonymous):

I'm not understanding what they are asking for.